# FPBoost: Fully Parametric Gradient Boosting for Survival Analysis

### v1

@@ Abstract
Survival analysis is a critical tool for analyzing time-to-event data and extracting valuable clinical insights. Recently, numerous machine learning techniques leveraging neural networks and decision trees have been developed for this task. Among these, the most successful approaches often rely on specific assumptions about the shape of the modeled hazard function. These assumptions include proportional hazard, accelerated failure time, or discrete estimation at a predefined set of time points. In this study, we propose a novel paradigm for survival model design based on the weighted sum of individual fully parametric hazard contributions. We build upon well-known ensemble techniques to deliver a novel contribution to the field by applying additive hazard functions, improving over approaches based on survival or cumulative hazard functions. Furthermore, the proposed model, which we call FPBoost, is the first algorithm to directly optimize the survival likelihood via gradient boosting. We evaluated our approach across a diverse set of datasets, comparing it against a variety of state-of-the-art models. The results demonstrate that FPBoost improves risk estimation, according to both concordance and calibration metrics.
@@ 1 Introduction
Survival analysis is a field of statistics with a pivotal role in healthcare data analysis, providing the ability to estimate the timing and associated uncertainty of clinical events. This capability is essential for aiding physicians in making safety-critical decisions based on data. Beyond healthcare, survival analysis has found applications in various fields, such as predicting equipment failures in industry or forecasting customer churn in relationship management. This widespread adoption underscores the importance of temporal risk estimation in diverse real-world scenarios [1].
@@
In practical applications, the event of interest can take several forms. In healthcare contexts, for example, it may denote patient mortality, disease recurrence, or hospital discharge. As another example, in customer relationship management it might represent a client’s initial purchase [2].
@@
Traditional survival models often rely on simplifying assumptions, such as risk proportion over between different subjects being constant over time [3]. While potentially beneficial in limited data scenarios, these simplifying assumptions can constrain model generalization [4]. Machine learning techniques have allowed to advanced these models by incorporating tree-based estimations and neural networks, significantly enhancing generalization capabilities and enabling the modeling of non-linear interactions within high-dimensional features. However, these advanced models still operate under certain constraints, such as time discretization in neural-based approaches [5] or accelerated failure time assumptions in gradient boosting methods [6].
@@
In this context, we introduce Fully Parametric Gradient Boosting (FPBoost), a novel architecture designed to model hazard functions through the composition of multiple fully parametric hazard functions. Hazard functions are related to survival functions as they measure the instantaneous risk of a subject experiencing the event of interest. FPBoost aims to combine the strengths of ensemble learning with tree-based gradient boosting [7], offering a flexible model with robust generalization capabilities and minimal assumptions. Modeling hazard as a weighted sum of multiple, fully-parametric functions, referred to as heads, allows FPBoost to be trained by maximizing the survival likelihood. This, in turn, removes the need for simplified expressions such as partial likelihood [3] or discrete losses [5]. Additionally, the continuous nature of the learned survival functions allows for a fine-grained estimation of the event distribution, without the need for interpolation techniques [8]. Finally, the use of gradient-boosted trees with tabular data, which is the most common data format in survival applications, can be competitive against neural network techniques [9], as highlighted by the empirical results presented in this study.
@@
We evaluate FPBoost in the right-censored, single-event setting, which is the most common application of survival analysis. The performance of FPBoost is benchmarked against state-of-the-art survival models, which include both tree-based and neural network-based models. Performance metrics include the concordance index [10] and the integrated Brier score [11]. Our experiments demonstrate that FPBoost outperforms alternative models in both concordance and calibration in the majority of cases and matches their performance when it does not.
@@
  • We present a detailed description of the FPBoost model, analyzing the composition of its hazard function, defining the survival likelihood for fully parametric heads, and outlining the training procedure based on gradient boosting.
  • We provide comprehensive empirical evidence of our approach’s robustness and efficacy through extensive experimentation across a diverse array of datasets and baseline models. The experimental procedure is designed to accommodate the intrinsic variability of survival datasets with low cardinality, providing an accurate estimation of the true generalization performance of each model under consideration.
  • We provide an open source Python implementation of FPBoost fully compatible with the scikit-survival library [12].
@@ 2 Background and Related Work
Survival analysis is a field of statistics focused on modeling the probability of an event of interest occurring over time for a population of subjects. The primary objective of survival models is thus to estimate a survival function $S(t|\mathbf{x})$, which measures the probability of surviving, i.e., not experiencing the event up to time $t$ as
@@
Here, $T$ is the time-to-event random variable and $\mathbf{x}\in\mathbb{R}^{d}$ a $d$-dimensional vector encoding the subject’s features. The survival function exhibits several key properties. It is monotonically non-increasing, starts at 1 for $t=0$, and asymptotically approaches 0 as $t$ tends to infinity indicating that, given an infinite time frame, all subjects will ultimately experience the event of interest [2].
@@ 2.1 Non-Parametric Models
Survival models can be cathegorized into three groups: non-parametric, semi-parametric, and fully parametric [2]. The group of non-parametric models comprises statistical estimators that provide information about data without any prior assumption on the event distribution. Non-parametric models rely on some notion of similarity between groups of individuals to improve the prediciton complexity. The most common non-parametric model is the Kaplan-Meier estimator [1], which is often used to plot the general survival behavior of a population. The KM estimator is defined as
$S(t)=\prod_{t_{i}\leq t}\left(1-\frac{e_{i}}{r_{i}}\right)$
where $t_{i}$ are the unique observed event times, $e_{i}$ is the number of events at time $t_{i}$, and $r_{i}$ is the number of subjects at risk before $t_{i}$. Notably, the KM estimator is not conditioned on the subjects’ features as it is tailored to provide aggregate information about the overall event distribution within the data. Another popular non-parametric model is the Random Survival Forest [13], which builds a set of decision trees with the CART [14] method by maximizing the event distribution difference between nodes. Each leaf contains a non-parametric estimation of the subjects corresponding to the leaf. The final prediction is obtained by averaging the predictions of the trees in the forest.
@@ 2.2 Semi-Parametric Models
Semi-parametric models are crucial tools in survival analysis. These models focus on predicting the hazard function, a quantity related to the survival function, which measures the instantaneous risk for subjects that have survived up to time $t$:
@@
The survival function is related to the hazard function as
@@
where $H(t|\mathbf{x})$ is the cumulative hazard function, defined as the integral of the hazard function from 0 to $t$ [2]. The Cox model [3] is a prominent semi-parametric model and serves as a primary baseline for machine learning-based survival analysis. This model relies on two key assumptions: (i) linear dependency between features and risk of experiencing an event (ii) the ratio between hazard functions of different subjects is constant over time. This latter is often referred to as proportional hazard assumption. While potentially limiting when the model is applied to large datasets, these assumptions provide a strong bias, enabling effective generalization even with limited data samples. In particular, the Cox model defines the hazard function as the product of a baseline hazard, $h_{0}(t)$, and the exponential of a subject-dependent risk factor $r(\mathbf{x})$:
@@
where $h_{0}(t)$ is a non-parametric hazard estimation common to all samples, such as the Breslow estimator [15]. The parameters $\beta$ are trained using the partial log-likelihood loss:
@@
Several extensions of the Cox model have been proposed, all relying on the proportional hazard assumption and partial log-likelihood optimization. Among those, CoxBoost [16] optimizes the partial log-likelihood using gradient boosting, leveraging decision trees to estimate $r(\mathbf{x})$. On the other hand, DeepSurv [4] replaces the linear dependency between parameters and features with a non-linear estimation using a single-output neural networks.
@@ 2.3 Fully Parametric Models
Building upon standard fully parametric distributions, Deep Survival Machines (DSM) [17] propose a parameter estimation neural network to construct a mixture of predefined probability distributions. The final survival function is then computed as a weighted sum of these distributions. DSM is trained using a combination of ELBO losses and a regularization prior loss in a Bayesian framework.
@@
Other popular models do not rely on predefined statistical distributions to construct survival estimations, but leverage neural networks to estimate the event probability directly at a fixed set of time intervals [18, 5]. These neural networks have a single output per time bin, representing the event probability for that interval. One such model is DeepHit [19], a discrete-time survival model capable of handling multiple event types. It consists of a shared feature extraction network followed by event-specific sub-networks estimating the probabilities for each event. While these discrete-time models have shown promising practical results, they may struggle with fine-grained or long-term prediction horizons due to their fixed-time nature. To address this limitation, some studies have proposed interpolation techniques between time points [5, 8].
3 Fully Parametric Survival Boosting
@@
Figure 1: The FPBoost architecture. In this example, a set of gradient-boosted trees estimate two distribution parameters, $\eta$ and $k$, for each of four heads. Heads 1 and 2 follow Weibull distributions, while heads 3 and 4 follow LogLogistic distributions. An additional set of trees estimates a positive weight for each head. These heads are combined to form a single hazard function and its corresponding cumulative hazard function. Trees are trained using the gradient of a loss combining negative log-likelihood and ElasticNet regularization.
@@
This work proposes FPBoost, a novel technique for survival analysis, based on the weighted sum of fully parametric hazard functions with trainable parameters. Parameter estimation is carried out via gradient boosting, optimizing the negative log-likelihood loss function.
The hazard function of FPBoost is composed of $J$ heads, each corresponding to a parametric distribution. In this work, heads follow either Weibull or LogLogistic distributions. However, any distribution differentiable with respect to its parameters can be used. Figure 1 depicts an example of a 4-headed FPBoost architecture with heads 1 and 2 following a Weibull distribution and heads 3 and 4 a LogLogistic distribution. As shown in Table 1, these distributions depend on two parameters: a scale parameter $\eta$ and a shape parameter $k$. The distribution parameters are estimated from the input features $\mathbf{x}$ using a set of regression trees. Additionally, another set of trees estimates a set of non-negative weighting parameters $w$, one for each head. The hazard function of FPBoost is defined as
$h(t|\Theta)=\sum\limits_{j=1}^{J}w_{j}h_{j}(t|\eta_{j},k_{j})$
where $w_{j}$ are the learned head weights and $\eta_{j},k_{j}$ are the parameter estimations from the regression trees for the $j$-th head. To improve readability, we define $\Theta$ as the vector containing all the distribution parameters $\eta_{j},k_{j}$ and head weighting factors $w_{j}$ for each of the $J$ heads. This way, each entry of $\Theta$ can be estimated with any regression model given input $\mathbf{x}$. To ensure that the distribution parameters and weighting factors are non-negative, we apply a ReLU activation function to $\Theta$, setting negative estimations to 0 [20].
Notably, the FPBoost hazard is differentiable with respect to the distribution parameters and weights. Therefore, FPBoost is trained by minimizing the negative log-likelihood loss function, which derives from equations 1 and 2:
$L_{\text{lik}}=-\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{i}\log(h(t_{i}|\Theta))-H(t_{i}|\Theta).$
In order to prevent overfitting, we add to $L_{\text{lik}}$ an ElasticNet regularization term [21] as

### v2

@@ Abstract
Survival analysis is a statistical framework for modeling time-to-event data. It plays a pivotal role in medicine, reliability engineering, and social science research, where understanding event dynamics even with few data samples is critical. Recent advancements in machine learning, particularly those employing neural networks and decision trees, have introduced sophisticated algorithms for survival modeling. However, many of these methods rely on restrictive assumptions about the underlying event-time distribution, such as proportional hazard, time discretization, or accelerated failure time. In this study, we propose FPBoost, a survival model that combines a weighted sum of fully parametric hazard functions with gradient boosting. Distribution parameters are estimated with decision trees trained by maximizing the full survival likelihood. We show how FPBoost is a universal approximator of hazard functions, offering full event-time modeling flexibility while maintaining interpretability through the use of well-established parametric distributions. We evaluate concordance and calibration of FPBoost across multiple benchmark datasets, showcasing its robustness and versatility as a new tool for survival estimation.
Machine Learning, Survival Analysis, Gradient Boosting, XGBoost, FPBoost
@@ 1 Introduction
Survival analysis is a field of statistics that plays a central role in data analysis for healthcare, providing the ability to estimate the timing and associated uncertainty of clinical events. This capability is essential to help physicians make informed safety-critical decisions based on data. Beyond healthcare, survival analysis has found applications in various fields, such as predicting equipment failures in industry or forecasting customer churn in relationship management. This widespread adoption underscores the importance of temporal risk estimation in various real-world scenarios (Wang et al., 2019).
@@
In practical applications, the event of interest can take several forms. In the healthcare context, for example, it can denote patient mortality, disease recurrence, or hospital discharge. As another example, in customer relationship management it might represent a client’s initial purchase (Klein et al., 2003).
@@
In order to model the survival function, traditional methods often rely on simplifying assumptions, such as the risk proportion between different subjects being constant over time (Cox, 1972). These assumptions allow the construction of survival functions from a small set of parameters that can be estimated with statistical methods. While suited for contexts with limited data, these simplifying assumptions constrain generalization in real-world scenarios (Katzman et al., 2018). Machine learning techniques have allowed to advance these models by incorporating decision trees and neural networks, significantly enhancing their ability to identify and learn non-linear interactions within high-dimensional features. However, most of these techniques still operate under certain constraints, such as time discretization in neural-based approaches (Kvamme & Borgan, 2021) or accelerated failure time in tree-based gradient boosting (Collett, 2023).
@@
In this context, we introduce Fully Parametric Gradient Boosting (FPBoost), a novel architecture designed to model hazard functions through the composition of multiple fully parametric hazard functions. Hazard functions are related to survival functions as they measure the instantaneous risk of a subject experiencing the event of interest. FPBoost combines the strengths of tree-based ensemble learning with gradient boosting (Friedman, 2001; Collett, 2023), offering a flexible model with robust generalization capabilities and minimal assumptions. Modeling hazard as a weighted sum of multiple, fully-parametric functions, referred to as heads, allows FPBoost to be trained by maximizing the full survival likelihood (Wang et al., 2019). This, in turn, removes the need for simplified assumptions such as partial likelihood (Cox, 1972; Katzman et al., 2018) or discrete losses (Kvamme & Borgan, 2021). Additionally, the continuous nature of the learned survival functions ensures a fine-grained estimation of the event distribution, without requiring interpolation techniques (Archetti et al., 2024). We theoretically show how this framework makes FPBoost a universal approximator of hazard functions, allowing it to learn in principle any target hazard, provided enough heads. Lastly, gradient-boosted trees applied to tabular data – the most common data format in survival applications – has proven to be still competitive against neural network techniques (Grinsztajn et al., 2022), as highlighted by our empirical results.
@@
We evaluate FPBoost in the right-censored, single-event setting, which is the most common application of survival analysis. The performance of FPBoost is benchmarked against state-of-the-art survival models, including both tree-based and neural network-based models (Chen et al., 2024). Performance metrics include the concordance index (Uno et al., 2011) to measure the discrimination capabilities of the model and the integrated Brier score (Graf et al., 1999), tailored for calibration. Our experiments demonstrate that FPBoost outperforms alternative models in both discrimination and calibration in the majority of cases and matches their performance when it does not.
@@
  • A detailed description of the FPBoost model, motivating the choices behind the composition of its hazard function, and detailing the training procedure based on gradient boosting.
  • A theoretical analysis on the approximation capabilities of FPBoost showing that it can model, in principle, any target hazard function, provided enough heads.
  • An extensive empirical analysis providing evidence of our approach’s efficacy across a diverse set of datasets and baseline models. The experimental procedure is designed to accommodate the intrinsic variability of survival datasets with low cardinality, providing an accurate estimation of the true generalization performance of each model under consideration.
  • An open-source Python implementation of FPBoost fully compatible with the scikit-survival library (Pölsterl, 2020). This way, the FPBoost algorithm can be directly used as a drop-in replacement in existing pipelines.
@@ 2 Background and Related Work
Survival analysis is a field of statistics that focuses on modeling the probability of an event of interest occurring over time for a population. The primary objective of survival models is thus estimating a survival function $S(t|\mathbf{x})$, which measures the probability of surviving, i.e., not experiencing the event up to time $t$ as
@@
Here, $T$ is the time-to-event random variable and $\mathbf{x}\in\mathbb{R}^{d}$ a $d$-dimensional vector encoding the subject’s features.
The survival function exhibits several key properties. It is monotonically non-increasing, starts at 1 for $t=0$, and asymptotically approaches 0 as $t$ tends to infinity indicating that, given an infinite time frame, all subjects will ultimately experience the event of interest (Klein et al., 2003).
@@ 2.1 Non-Parametric Models
Survival models can be categorized into three groups: non-parametric, semi-parametric, and fully parametric (Wang et al., 2019). The group of non-parametric models comprises statistical estimators that provide information about data without any prior assumption on the event distribution. Non-parametric models rely on some notion of similarity between groups of individuals to improve the prediciton complexity. The most common non-parametric model is the Kaplan-Meier (KM) estimator (Goel et al., 2010), which is often used to plot the general survival behavior of a population. In fact, the KM estimator is not conditioned on the subjects’ features as it is tailored to provide aggregate information about the overall event distribution within the data. Another popular non-parametric model is the Random Survival Forest (Ishwaran et al., 2008), which builds a set of decision trees with the CART (Breiman et al., 1984) method by maximizing the event distribution difference between nodes according to repeated log-rank tests (Bland & Altman, 2004). Each leaf contains a non-parametric estimation of the subjects corresponding assigned to that specific terminal node. The final prediction is obtained by averaging the predictions of the trees in the forest.
@@ 2.2 Semi-Parametric Models
Semi-parametric models are crucial tools for survival analysis, providing the ability to build survival estimations from a combination of non-parametric and parametric techniques. These models focus on predicting the hazard function, a quantity related to the survival function, which measures the instantaneous risk for subjects that have survived up to time $t$:
@@
Differently from the survival function, which is constrained between 0 and 1, the hazard function can take values greater than 1. Additionaly, the survival function is related to the hazard function as
@@
where $H(t|\mathbf{x})$ is the cumulative hazard function, defined as the integral of the hazard function from 0 to $t$. For this reason, the cumulative hazard diverges for $t\rightarrow\infty$ to allow the survival function to asymptotically approach 0 (Wang et al., 2019).
The Cox model (Cox, 1972) is a prominent semi-parametric model and serves as a primary baseline for machine learning-based survival analysis. This model relies on two key assumptions: (i) linear dependency between features and risk of experiencing an event (ii) the ratio between hazard functions of different subjects is constant over time. This latter is often referred to as proportional hazard assumption. While potentially limiting when the model is applied to large datasets, these assumptions provide a strong bias, enabling effective generalization even with limited data samples. In particular, the Cox model defines the hazard function as the product of a baseline hazard, $h_{0}(t)$, and the exponential of a subject-dependent risk factor $r(\mathbf{x})$:
@@
where $h_{0}(t)$ is a non-parametric hazard estimation common to all samples, such as the Breslow estimator (Breslow, 1974). The parameters $\beta$ are trained using the partial log-likelihood loss:
@@
Several extensions of the Cox model have been proposed, all relying on the proportional hazard assumption and partial log-likelihood optimization. Among those, CoxBoost (Ridgeway, 1999) and XGBoost (Chen & Guestrin, 2016) optimize the loss using gradient-boosted decision trees to estimate $r(\mathbf{x})$. On the other hand, DeepSurv (Katzman et al., 2018) replaces the linear dependency between parameters and features with a single-output neural networks.
@@ 2.3 Fully Parametric Models
Building upon standard fully parametric distributions, Deep Survival Machines (DSM) (Nagpal et al., 2021) propose a parameter estimation neural network to construct a mixture of predefined probability distributions. The final survival function is then computed as a weighted sum of these distributions. DSM is trained using a combination of ELBO losses and a regularization prior loss in a Bayesian framework.
@@
Other popular models do not rely on predefined statistical distributions to construct survival estimations, but leverage neural networks to estimate the event probability directly at a fixed set of time intervals (Kvamme et al., 2019; Kvamme & Borgan, 2021). These neural networks have a single output per time bin, representing the event probability for that interval. One such model is DeepHit (Lee et al., 2018), a discrete-time survival model consisting of a shared feature extractor followed by event-specific sub-networks estimating the probabilities for each event. While these discrete-time models have shown promising practical results, they struggle with fine-grained or long-term prediction horizons due to their fixed-time nature. To address this limitation, some studies have proposed interpolation techniques between time points (Kvamme & Borgan, 2021; Archetti et al., 2024).
3 Fully Parametric Gradient Boosting
@@
Figure 1: FPBoost architecture example with four heads. A set of trees estimates two distribution parameters, $\eta_{j}$ and $k_{j}$, for each of four heads starting from the input features. Heads 1 (blue) and 2 (green) follow Weibull distributions, while heads 3 (orange) and 4 (yellow) follow LogLogistic distributions. An additional set of trees (gray) estimates a weight for each head. These heads are combined to form a single hazard function and its corresponding cumulative hazard function. New trees are built by fitting the gradient of the negative log-likelihood and ElasticNet (purple).
FPBoost is a novel survival algorithm based on the weighted sum of fully parametric hazard functions. Parameter estimation is carried out via gradient boosting, optimizing the negative log-likelihood loss function.
3.1 Hazard Function Definition
In FPBoost, the hazard function is composed of $J$ heads, each corresponding to a parametric distribution. This work includes heads following either the Weibull or LogLogistic distributions but, in principle, any hazard distribution differentiable with respect to its parameters can be included as a head. Table 1 collects the formulation of the considered distributions. This choice comes primarily from the fact that Weibull distributions are suited to model constant and increasing risks, while LogLogistic ones provide decreasing and arc-shaped risk profiles. Typically, the former is associated to survival behaviors such as aging and wear, while the latter relates to infant mortality. When combined, the resulting risk can assume a bathtub shape, covering a wide range of real-world behaviors (Klein et al., 2003; Nagpal et al., 2021). This approach provides an advantage from both a learning and interpretation perspective, as each sample is associated with a set of well-known distribution functions, where each parameter has a clear interpretation. In particular, for both the considered distributions, the $\eta$ parameter controls the scale – or spread – while the $k$ parameter changes the curve shape by modifying the slope or adding a bump.
Figure 1 depicts an example of a 4-headed FPBoost architecture with heads 1 and 2 following a Weibull distribution and heads 3 and 4 a LogLogistic distribution. The distribution parameters are estimated from the input features $\mathbf{x}$ using a set of regression trees. Additionally, another set of trees estimates a weight $w_{j}$ for each head. Thus, the hazard function of FPBoost is defined as
$h(t|\Theta)=\sum\limits_{j=1}^{J}w_{j}h_{j}(t|\eta_{j},k_{j})$
where $w_{j}$ are the learned head weights and $\eta_{j},k_{j}$ are the parameter estimations from the regression trees for the $j$-th head. To improve readability, we define $\Theta$ as the vector containing all the distribution parameters $\eta_{j},k_{j}$ and head weighting factors $w_{j}$ for each of the $J$ heads.
@@ 3.1 Hazard Function Definition
To guarantee the validity of the Weibull and LogLogistic hazard formulations, the scale and shape parameters $\eta_{j},k_{j}$ must be nonnegative. To enforce this constraint, we apply a ReLU activation function (Nair & Hinton, 2010) to these parameters, ensuring that any negative estimates are set to zero.
Similarly, the weighting parameters $w_{j}$ are processed by an activation function $\phi(\cdot)$ before hazard computation. The choice of activation function significantly influences model interpretability and generalizability. In particular, functions that yield nonnegative values, such as ReLU, sigmoid, and softmax, promote interpretability, as they quantify the contribution of each hazard component with a positive weight. This approach allows for meaningful insights into the relative contributions of Weibull and LogLogistic heads in hazard estimation. For instance, a predominance of Weibull-based hazards could indicate a higher susceptibility to aging-related failures, whereas a stronger presence of LogLogistic hazards might suggest risks associated with early failure or infant mortality.
The use of activation functions $\phi(\cdot)$ that allow for negative weights, despite making interpretation less straightforward, enables a a broader modeling capacity, as demonstrated by the following result:
Theorem 3.1.
[…]
$\sup_{t\in[0,T]}\biggl{|}\,h^{\star}(t)\;-\;\sum\limits_{j=1}^{J}w_{j}\,h^{W}_{j}(t)\biggr{|}\;<\;\varepsilon.$
The proof, detailed in Appendix A.1, is a direct consequence of the Weierstrass Approximation Theorem, after showing that a single Weibull head is equivalent to a monomial of arbitrary degree. This underpins the capability of FPBoost to approximate any target hazard function with sufficient heads. To the best of our knowledge, this represents the first universal approximation theorem for machine learning models in the context of survival analysis. It is important to note that while only Weibull hazards are required for this theoretical result, incorporating LogLogistic hazards often enhances model performance and makes optimization easier in practical applications.
On a final technical note, allowing negative weights means the final hazard function could potentially be negative for some values of $t$. Although this does not compromise the theoretical result, the FPBoost implementation prevents this issue by clipping the final hazard value to be above zero.
3.2 Loss Function and Training Procedure
Since the FPBoost hazard is differentiable with respect to the distribution parameters and weights, the estimators can be trained by minimizing the full negative log-likelihood loss function, which derives from equations 1 and 2, without the need of simplifying assumptions:
$L_{\text{lik}}=-\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{i}\log(h(t_{i}|\Theta_{i}))-H(t_{i}|\Theta_{i}).$
In order to prevent overfitting, we add to $L_{\text{lik}}$ an ElasticNet regularization term (Zou & Hastie, 2005) as